### v2

# DiCeption: A Generalist Diffusion Model for Visual Perceptual Tasks

Abstract
Our primary goal here is to create a good, generalist perception model that can tackle multiple tasks, within limits on computational resources and training data. To achieve this, we resort to text-to-image diffusion models pre-trained on billions of images and introduce our visual generalist model: DiCeption. Our exhaustive evaluation metrics demonstrate that DiCeption effectively tackles multiple perception tasks, achieving performance on par with state-of-the-art models. We achieve results on par with SAM-vit-h using only 0.06% of their data (e.g., 600K vs. 1B pixel-level annotated images). Inspired by Wang et al. [109], DiCeption formulates the outputs of various perception tasks using color encoding; and we show that the strategy of assigning random colors to different instances is highly effective in both entity segmentation and semantic segmentation. Unifying various perception tasks as conditional image generation enables us to fully leverage pre-trained text-to-image models. Thus, DiCeption can be efficiently trained at a cost of orders of magnitude lower, compared to conventional models that were trained from scratch. When adapting our model to other tasks, it only requires fine-tuning on as few as 50 images and $\sim$1of its parameters. DiCeption provides valuable insights and a more promising solution for visual generalist models. Project webpage, and huggingface demo are available.
1 Introduction
Figure 1: With one single model, DiCeption solves multiple tasks without relying on any task-specific modules (rows 1 to 3). The red dots in the figure indicate the input points used for point-prompted segmentation. DiCeption preserves fine details in segmentation, such as hair (row 4). DiCeption supports both human pose estimation and semantic segmentation (row 5, 6). DiCeption can quickly adapt to new tasks by fine-tuning less than 1% of its parameters on as few as 50 images (row 7). For additional visualizations, please refer to Figures S1, S2, S4, S6, S7, S8, S9, S10, S11, S12, S13, S14 in the Appendix.
Foundation models [49, 87, 118, 119, 117, 11, 7, 83, 75, 91, 6, 39], typically requiring extensive training on billions of data samples, play a pivotal role in their respective domains. In natural language processing (NLP), current foundation models [9, 99, 100, 26] have already demonstrated the potential to serve as versatile solutions, solving diverse fundamental tasks and with minimal fine-tuning needed for new tasks. This success can be attributed to the relatively small representational differences among various language tasks. However, in the domain of computer vision, task representations can differ substantially, and up to date, we still lack an effective approach to unify these distinct tasks. Consequently, existing vision foundation models usually excel at one single specific task, such as image segmentation [49, 87] or monocular depth estimation [118, 119, 117], because they are trained on data tailored exclusively to that task. Owing to the pronounced disparity in visual representations across tasks, coupled with the single-task specialization that characterizes current vision foundation models, fine-tuning these models for new tasks remains a formidable challenge. Although some efforts [12, 75, 39, 88] have been made to learn universal visual representations for more generalized vision foundation models, their performance still falls noticeably short compared to specialized models.
Recent studies [109, 69, 68, 72, 2, 113] on visual generalist models are predominantly trained from scratch, often requiring substantial computational resources and large datasets to achieve good results. Unfortunately, the price of collecting a sufficiently large and high-quality multi-task dataset is substantial. Here, inspired by the success of diffusion models, we propose the hypothesis that leveraging their powerful priors can help mitigate the significant computational and data overhead for training powerful generalist models. While some existing works [47, 114, 38, 123, 93] have demonstrated that this is feasible in single-task scenarios, the potential of diffusion model priors in multi-task settings remains largely under-explored.
In this paper, we successfully leverage the priors of diffusion models to achieve results on par or even better with the state-of-the-art models on various tasks with only minimal training data. We name our powerful visual generalist model DiCeption. For each task, we require substantially less data than specialized foundation models. For instance, compared to SAM segmentation, DiCeption  achieves comparable performance using a significantly smaller dataset of 600K samples, without any selective sampling. This is in sharp contrast to SAM which relies on a massive dataset of 1 billion pixel-level annotated samples. This demonstrates the efficiency and robustness of DiCeption, as it requires only 0.06% of the data used by SAM while still producing competitive results.
The concurrent work One Diffusion [53] proposes a unified model for image generation by treating different images as different views. However, their approach fails to distinguish between conditions and images, greatly limiting its effectiveness in perception tasks. More critically, their from-scratch training demands substantial data and computational resources. In contrast, our work focuses on image perception rather than generation, achieving on-par state-of-the-art performance in perception tasks. See more details in Appendix 2.4. DiCeption  highlights that the generative image priors lead to surprisingly more efficient and effective pathways to generalist image understanding models. The entire training process is remarkably stable, eliminating the need to design a complex training recipe. Even more notably, DiCeption is capable of quickly adapting to new tasks using as few as 50 training images and fine-tuning less than 1% of its parameters. Moreover, we show that previously ineffective strategies in generalist models such as task control via prompts [113] and random color assignments for segmentation masks [109] are viable in our framework. Additionally, we investigate the role of 1-step denoising in multi-task settings, a technique proven beneficial in single-task settings. DiCeption provides valuable insights for the design and training of strong visual generalist models.
In summary, our main contributions are as follows.
  • We introduce DiCeption, a generalist model capable of performing multiple visual perception tasks. Extensive experiments demonstrate that DiCeption can effectively address various tasks, achieving or par accuracy with state-of-the-art specialist models.
  • When trained on new tasks, DiCeption achieves high-quality results with minimal fine-tuning. Using only a small number of images (even with 50 images), the model can achieve exceptional performance by fine-tuning less than 1% of its parameters.
  • We conduct a thorough set of experiments exploring various aspects of multitask models, including network architectures, and 1-step diffusion. We believe that our experimental findings offer valuable insights and will significantly benefit future research.
2 Related Work
2.1 Vision Foundation Models
Vision foundation models are models that are trained on large-scale datasets and demonstrate excellent performance within their trained domains. Vision foundation models now exist for a broad range of vision tasks, including monocular depth estimation [118, 119, 117, 7], object detection [11], segmentation [49, 87], multimodal tasks [83, 64], image and video generation [91, 28, 6], and more recently, emerging 3D models [105, 70]. While many works [110, 48, 58, 84, 133, 136] have sought to leverage the prior knowledge embedded in these models to tackle other tasks, such efforts often require complex network designs and intricate training strategies, typically transferring only to a limited number of tasks. Some foundation models [88, 39, 75, 12] emphasize representation learning, aiming to solve diverse downstream tasks by relying on generalized features. However, the results of these methods often fall short when compared with specialized foundation models. In contrast, our approach ensures consistent accuracy across multiple tasks while also enabling swift adaptation to new downstream tasks.
2.2 Diffusion Models
Diffusion models [28, 91, 6] have achieved remarkable success in image and video generation in recent years. The idea is to gradually add noise to the data and train a model to reverse this process, denoising step by step to generate the result. Recent diffusion models [28] utilize flow matching [63, 1, 66] and the DiT architecture [77], making them more scalable and efficient. Diffusion models have enabled a wide range of notable applications, including conditional image generation [130, 61, 124, 73, 82], image editing [8, 46, 116], story generation [107, 135], video generation [40, 35, 132, 122, 6, 50, 106], and video editing [13, 65, 14]. These successes underscore the substantial prior knowledge embedded in diffusion models.
Building on this insight, many studies [114, 38, 123, 47, 136] leverage diffusion models for downstream image understanding tasks. However, these approaches typically require separate fine-tuning for each individual task. Recently, we find several concurrent works [112, 53] also use diffusion models for multitask learning. Yet, these methods often involve complex network architectures and training procedures, and their evaluations tend to focus only on a very limited subset of image understanding results. In contrast, our DiCeption offers a simpler solution. We not only conduct detailed evaluations of our method across a variety of tasks but also demonstrate that the simplicity, paired with the inherent strengths of diffusion models, can be sufficient to deliver strong results without relying on overly complicated setups.
2.3 Multi-task Generalist Models
Recently, there has been a surge of interest in exploring visual multitask learning. Some approaches [109, 111] draw inspiration from in-context learning in NLP, adapting it for the visual domain. Others [69, 68, 72, 2] have advocated for sequence modeling methods, utilizing a transformer encoder-decoder architecture. In these approaches, different encoders map various tasks into a shared representation space, and distinct decoders are employed to transform tokens into the outputs specific to each task. However, these methods face notable limitations: they need to train a separate encoder and decoder for every individual task and they usually rely on substantial amounts of data to attain optimal performance.
The recent success of high-quality Vision Language Models (VLMs) [64] has also encouraged researchers to leverage them for building multitask models. Yet, these VLM-based methods [4, 104, 16, 67, 89, 59] typically focus on multimodal understanding tasks, such as image captioning, rather than general visual perception tasks. Meanwhile, some approaches [97, 132, 76] combine diffusion models with autoregressive models, focusing primarily on instruction-following image generation or editing tasks, rather than addressing image perception tasks. Although certain studies [52, 45, 17, 34] have tried to apply VLMs to more advanced semantic perception tasks, they struggle to establish a unified generalist visual model.
2.4 Compared with One Diffusion
The concurrent work, One Diffusion [53], addresses multi-task image generation, whereas our approach focuses on multi-task image understanding. We excel at performing a broader range of image understanding tasks with higher quality. While One Diffusion’s strategy of treating different images as different views benefits generation tasks, their failure to distinguish between conditions and images introduces harmful degrees of freedom for perception tasks, as illustrated in the red-highlighted regions of Figure S5. Specifically, when performing perception tasks, One Diffusion tends to generate an image similar to the original input, rather than the desired perceptual results.
Although One Diffusion suggests that more detailed text prompts can lead to better results, we argue that performance in perception tasks should not overly depend on the quality of text prompts. In contrast, our method uses only simple task prompts to distinguish between different tasks, rather than allowing the text prompts to dominate the results.
Crucially, while One Diffusion requires a massive amount of data (75 million samples) and computational resources for from-scratch training, we leverage the priors of pretrained models and demonstrate that, with significantly less data (1.8 million samples), we achieve performance on par with state-of-the-art results. In the image understanding tasks shared by both approaches, we consistently produce more stable and higher-quality results than One Diffusion.
3 Method
3.1 Preliminary
Recent diffusion models predominantly build upon flow matching methodologies [63, 1, 66] and the DiT architecture [77]. They aim to learn a velocity field that effectively maps samples from a source distribution to a target distribution. The training process involves minimizing the discrepancy between the model’s predicted velocity and the ground-truth velocity, which is typically formulated as:
${\rm Loss}={\mathbb{E}}_{\mathbf{z}_{0},t,c}\|v_{\theta}(\mathbf{z}_{t},t,c)-u(\mathbf{z}_{t})\|^{2}_{2},$ (1)
where $c$ is the condition, usually the text prompt. $\mathbf{z}_{0}$ is the clean image latent and $\mathbf{z}_{t}$ is the noisy one. The learned velocity field corresponds to an ordinary differential equation (ODE), such that during inference, a sample drawn from the source distribution can be transformed into the desired output by solving this ODE.
3.2 Unifying Task Representation into RGB Space
The decision to unify representations of diverse tasks in RGB space was motivated by two key factors: (1) It Maximally leverages the priors in text-to-image models, which have been extensively trained within the RGB domain. (2) RGB serves as a foundational representation in computer vision, providing a common visual framework through which a wide variety of tasks can be coherently and intuitively visualized.
We focus on several of the most fundamental tasks in computer vision: monocular depth estimation, normal estimation, and segmentation. Segmentation, in particular, encompasses point-prompted segmentation, entity segmentation, and semantic segmentation. All these tasks can be unified within an RGB space, with the difference being the number of channels. For single-channel representations, such as depth maps and segmentation masks, we align them with RGB by repeating the channel data three times. For inherently three-channel representations, such as normal maps, we treat them directly as RGB images.
Entity segmentation is to segment every instance in an image but with no category. We assign each mask within an image a random color and merge them into a three-channel RGB mask. Painter found that assigning color randomly makes the model hard to optimize. However, we find this approach has no adverse impact on the training and enables the model to effectively learn to distinguish different instances by painting them with different colors. Each instance’s mask can be extracted from the RGB mask using clustering algorithms during post-processing without significant performance degradation. We also apply the random color assignment in semantic segmentation. Unlike traditional semantic segmentation, our method is capable of segmenting instances of the same semantic category. By default, We use KMeans for mask extraction.
Let $\mathbf{x}_{r}$ denote the pre-unified raw representation for each task, and $\mathbf{x}$ represents the unified RGB-like output representation. We formalize this process as: $\mathbf{x}=\Psi(\mathbf{x}_{r})$.
3.3 DiCeption: A Unified Framework
Architecture.
Our model adopts the same architecture as SD3. We aim to keep the architecture as unchanged as possible, fully leveraging the pretrained prior knowledge. We use simple task prompts to direct the model to perform various tasks, such as “image2depth,” “image2normal,” and “image2segmentation.”
For point-prompted segmentation, a naive approach is directly painting points on the image. But this strategy is highly sensitive to the size of the points. If the painted points are too large, they can obscure small regions, causing segmentation to fail. Conversely, if the painted points are too small, the model may lose relevant point information after VAE downsampling and patchification. To address this, we introduce a minimal straightforward two-layer MLP $\Phi(\cdot)$ that enables the model to understand the point prompt.
Inspired by SAM, we apply sin-cos positional encoding to the point coordinates $p$, then pass them into the MLP $\Phi(\cdot)$ to produce point embeddings that match the dimension of the input hidden states. We use two learnable embeddings to indicate whether the embedding is a point embedding or not: $\xi_{p}$ for point embeddings and $\xi_{np}$ for non-point embeddings. The processed point embedding is summed with $\xi_{p}$. For other tasks, we simply use $\xi_{np}$ as the point embedding. During training, we randomly select 1–5 points to guide the segmentation. When the number of selected points is fewer than 5, we pad the point embeddings to a length of 5 with $\xi_{np}$. When performing tasks that do not require point input, the point embedding is simply a length-5 sequence, where each element is $\xi_{np}$. By denoting the final point embedding as $\xi$, this process is formulated as:
$\xi=\begin{cases}\operatorname{Concat}(\Phi(\operatorname{PE}(p))+\xi_{p},\xi_{np})&\text{if point segmentation}\\ \xi_{np}&\text{else}\end{cases}$ (2)
Input Formulation and Loss.
DiCeption introduces two additional inputs based on SD3: the input image $\mathbf{x}^{\prime}$ and point embedding $\xi$. For the input image, we first apply VAE to down-sample it by a factor of 8, after which it is $2\times 2$ patchified into sequences. We denote this pre-processing as $\tau$. Subsequently, the task prompt token $\mathbf{e}$, point embedding $\xi$, noisy token $\mathbf{z}_{t}$, and input image token $\mathbf{z}^{\prime}$ are concatenated to form the complete input. We follow the flow matching loss in training SD3. During training, the loss is applied solely to the noisy tokens:
$\mathbf{z}_{0}$ $=\tau(\mathbf{x})$ (3)
$\mathbf{z}^{\prime}$ $=\tau(\mathbf{x}^{\prime})$
${\rm Loss}={\mathbb{E}}_{\mathbf{z}_{0},t}\|v_{\theta}$ $(\mathbf{z}_{t},\mathbf{z}^{\prime},t,\mathbf{e},\xi)-u(\mathbf{z}_{t})\|^{2}_{2}.$
3.4 Adapting to New Tasks
In practical scenarios, it is often necessary to enable models to quickly adapt to new tasks using only a small amount of training data. Traditional specialized foundation models, however, are typically limited to tasks closely related to their domain and generally require substantial datasets and carefully designed network architectures for adaptation. Diffusion models, while powerful in many respects, cannot easily adapt to downstream tasks through fine-tuning only a few parameters with limited data.
DiCeption effectively addresses this limitation. We conducted experiments on lung segmentation, tumor segmentation, and image highlighting, which represent tasks with varying degrees of overlap with the model’s original domain. We train fewer than 1% of the model’s parameters using LoRA [42] without any complex architectural modifications. Notably, despite the limited availability of training samples (50 per task), DiCeption consistently delivered successful and high-quality performance across all target tasks. These results provide compelling evidence for the potential of DiCeption as a truly unified foundation model.
4 Experiments
4.1 Implementation Details
Data.
We randomly select 500k images from the OpenImages [51] dataset and use DepthPro [7] and StableNormal [123] to generate depth and normal data. For point-prompted segmentation, we randomly select 400k images from the SA-1B [49] dataset, as well as 200k images with fine-grained hair masks synthesized from the AM2k [56], AIM500 [57], and P3M-10k [55]. Entity segmentation data is from EntityV2 [81], while semantic segmentation data comes from the COCO-Rem [94], and human pose data is sourced from COCO [62]. For few-shot fine-tuning, we select 50 samples from the publicly labeled Chest X-Ray dataset [108], LOL-v2 [121], and Kaggle’s Brain Tumor dataset as training samples. More details can be found in Appendix  S1.
Training.
[…]
Inference.
We follow the settings of the pretrained model and perform 28 steps of denoising during inference. The inference can be run on a GPU of 24GB memory with a batch size of 4.
4.2 Comparisons with Existing Methods
Table 1: Quantitative comparison of depth estimation with both specialized models and multi-task models on zero-shot datasets. Our visual generalist model can perform on par with state-of-the-art models. We use the same evaluation protocal ($\dagger$) as Genpercept [114].
Method Training KITTI [32] NYUv2 [74] ScanNet [23] DIODE [102] ETH3D [92]
[…]
MiDaS [86] 2M 0.236 0.630 0.111 0.885 0.121 0.846 0.332 0.715 0.184 0.752
Omnidata [27] 12.2M 0.149 0.835 0.074 0.945 0.075 0.936 0.339 0.742 0.166 0.778
DPT-large [85] 1.4M 0.100 0.901 0.098 0.903 0.082 0.934 0.182 0.758 0.078 0.946
DepthAnything [118] 63.5M 0.080 0.946 0.043 0.980 0.043 0.981 0.261 0.759 0.058 0.984
DepthAnything v2 [119] 62.6M 0.080 0.943 0.043 0.979 0.042 0.979 0.321 0.758 0.066 0.983
Depth Pro [7] - 0.055 0.974 0.042 0.977 0.041 0.978 0.217 0.764 0.043 0.974
Metric3D v2 [43] 16M 0.052 0.979 0.039 0.979 0.023 0.989 0.147 0.892 0.040 0.983
DiverseDepth [125] 320K 0.190 0.704 0.117 0.875 0.109 0.882 0.376 0.631 0.228 0.694
LeReS [126] 354K 0.149 0.784 0.090 0.916 0.091 0.917 0.271 0.766 0.171 0.777
HDN [128] 300K 0.115 0.867 0.069 0.948 0.080 0.939 0.246 0.780 0.121 0.833
GeoWizard [31] 280K 0.097 0.921 0.052 0.966 0.061 0.953 0.297 0.792 0.064 0.961
DepthFM [33] 63K 0.083 0.934 0.065 0.956 - - 0.225 0.800 - -
Marigold [47] 74K 0.099 0.916 0.055 0.964 0.064 0.951 0.308 0.773 0.065 0.960
DMP Official [54] - 0.240 0.622 0.109 0.891 0.146 0.814 0.361 0.706 0.128 0.857
GeoWizard [31] 280K 0.129 0.851 0.059 0.959 0.066 0.953 0.328 0.753 0.077 0.940
DepthFM [33] 63K 0.174 0.718 0.082 0.932 0.095 0.903 0.334 0.729 0.101 0.902
Genpercept [114] 90K 0.094 0.923 0.091 0.932 0.056 0.965 0.302 0.767 0.066 0.957
Painter [109] 24K 0.324 0.393 0.046 0.979 0.083 0.927 0.342 0.534 0.203 0.644
Unified-IO [69] 48K 0.188 0.699 0.059 0.970 0.063 0.965 0.369 0.906 0.103 0.906
4M-XL [72] 759M 0.105 0.896 0.068 0.951 0.065 0.955 0.331 0.734 0.070 0.953
OneDiffusion  [53] 500K 0.101 0.908 0.087 0.924 0.094 0.906 0.399 0.661 0.072 0.949
Ours-single 500K 0.081 0.942 0.068 0.949 0.078 0.945 0.267 0.709 0.059 0.969
Ours 500K 0.075 0.945 0.072 0.939 0.075 0.938 0.243 0.741 0.053 0.967
Table 2: Quantitative comparison of surface normal estimation with both specialized models and multi-task models. All methods are evaluated with the same evaluation protocol of StableNormal [123].
Method Training NYUv2 [74] ScanNet [23] DIODE-indoor [102]
Samples $\downarrow$ $\downarrow$ $11.25^{\circ}$$\uparrow$ $22.5^{\circ}$$\uparrow$ $30^{\circ}$$\uparrow$ $\downarrow$ $\downarrow$ $11.25^{\circ}$$\uparrow$ $22.5^{\circ}$$\uparrow$ $30^{\circ}$$\uparrow$ $\downarrow$ $\downarrow$ $11.25^{\circ}$$\uparrow$ $22.5^{\circ}$$\uparrow$ $30^{\circ}$$\uparrow$
DINSE [3] 160K 18.572 10.845 54.732 74.146 80.256 18.610 9.885 56.132 76.944 82.606 18.453 13.871 36.274 77.527 86.976
Geowizard [31] 280K 20.363 11.898 46.954 73.787 80.804 19.748 9.702 58.427 77.616 81.575 19.371 15.408 30.551 75.426 86.357
GenPercept [114] 90K 20.896 11.516 50.712 73.037 79.216 18.600 8.293 64.697 79.329 82.978 18.348 13.367 39.178 79.819 88.551
Marigold [47] 90K 20.864 11.134 50.457 73.003 79.332 18.463 8.442 64.727 79.559 83.199 16.671 12.084 45.776 82.076 89.879
StableNormal [123] 250K 19.707 10.527 53.042 75.889 81.723 17.248 8.057 66.655 81.134 84.632 13.701 9.460 63.447 86.309 92.107
Unified-IO [68] 210K 28.547 14.637 39.907 63.912 71.240 17.955 10.269 54.120 77.617 83.728 31.576 16.615 27.855 64.973 73.445
4M-XL [72] 759M 37.278 13.661 44.660 60.553 65.327 30.700 11.614 48.743 68.867 73.623 18.189 12.979 36.622 81.844 87.050
Ours-single 500K 18.267 10.238 52.393 76.802 83.113 19.892 12.424 45.930 74.341 81.965 17.611 8.912 62.030 80.827 86.474
Ours 500K 18.302 10.538 52.533 75.977 82.573 19.348 12.129 46.410 74.805 82.176 17.946 8.686 62.641 81.152 85.398
Table 3: Average recall (AR) of entity segmentation on the MS COCO validation set.
Method AR-$\uparrow$ AR-$\uparrow$ AR-$\uparrow$
EntityV2 [81] 0.313 0.551 0.683
Ours-single 0.123 0.424 0.648
Ours 0.121 0.439 0.637
Table 4: Evaluation of human keypoints estimation on MS COCO.
HRNet[96] HRFormer[129] ViTPose[115] Painter[109] Ours
[…]
Table 5: Evaluation of semantic segmentation on the MS COCO (with category ID).
Method SparK [98] OneFormer [44] Mask2Former [18] Ours
[…]
Figure 2: Comparisons of mIoU with SAM-vit-h. We achieve results on par with SAM using only 0.06% of their data (600K vs.  1B).
We compare the performance of specialized models, existing multi-task models, and our DiCeption across various tasks. Specifically, we evaluate depth using the same protocol as Genpercept, normal estimation using the same method as StableNormal, point-prompted segmentation using the same approach as SAM, and human keypoints using the same method as Painter. We also assess semantic and entity segmentation on the MS COCO dataset. For entity segmentation, we assigned all predicted categories to the same label.
As in Tables 1 and 2, our DiCeption significantly outperforms existing multi-task models and achieves performance on par with state-of-the-art specialized models or demonstrates only an acceptable performance decrease. For point-prompted segmentation, as shown in Figure 2, we achieve results on par with SAM-vit-h using only 0.06% of their data. SAM shows a clear advantage only on certain out-of-distribution datasets that are outside the scope of our model’s training, such as WoodScape fisheye dataset. Notably, while most specialized models require extensive data or complex data pipelines, our method achieves excellent results with significantly less data, the majority of which is obtained through simple random selection. For entity segmentation, we observe that our model performs poorly on small objects. We believe that this is due to the limited amount of data, which causes the model to overfit larger objects. This problem can be solved by introducing more data towards small objects.
We observe that, although our model generates high-quality visualizations for human pose and semantic segmentation, the corresponding evaluation metrics remain relatively low. For human keypoints, this is primarily due to two factors: Firstly, we utilize skeletal-form RGB images rather than heatmaps. While the former produces visually appealing results, the extraction of keypoints during post-processing introduces considerable errors. Secondly, our evaluation follows the 192$\times$256 top-down human keypoints protocol. The original 192$\times$256 images are resized to 768$\times$768 before being input into the model, resulting in extremely blurred inputs that likely contribute to the diminished performance.
Semantic segmentation also introduces considerable errors during post-processing. We offer a comprehensive explanation of the metric degradation in the Appendix B.3. For individual images, we can adjust hyperparameters during the post-processing of RGB mask to achieve the optimal masks. However, different results require different optimal hyperparameters, and manually adjusting each sample in the validation set during evaluation is impractical due to the excessive labor involved. This leads to remarkable semantic segmentation visualizations produced by our method but relatively low metrics. We achieve excellent results by training solely on COCO-Rem. Furthermore, our semantic segmentation can distinguish different instances of the same semantic category. We believe we have strongly demonstrated the potential of our method for instance-level semantic segmentation with IDs. Please check Appendix B for more details about post-processing.
4.3 Comparisons with Our Single-task Models
For the training of single-task models, we ensure that the network architecture remains the same and the total amount of training data seen for each specific task is the same as that for the multi-task model. For example, if the multi-task model trains for 100 iterations with 4 depth data samples per batch, the single-task model will also be trained for 100 iterations with 4 data samples per batch. In our current data setting (approximately 1.8 million samples), we have not observed a significant gap between the multi-task and single-task models, nor have we seen a trend of mutual promotion between different tasks, as shown by “Ours-single” in Tables 1, 2, 3 and Figure 2.
We believe that it is more appropriate to explore with larger datasets in order to draw more solid conclusions. We leave this as future work.
4.4 Ablations
Multi-point Prompted Segmentation.
Ambiguity is a significant issue in point-prompted segmentation. For example, if a point is placed on a person’s clothing, the model may segment the clothing, but the desired result is the person. Therefore, more points are needed to resolve this ambiguity. As illustrated in Figure 3, additional points help the model better segment the desired results.
Figure 3: Comparisons between 1-point and 5-point segmentation of mIoU on all 23 validation datasets.
Architecture of Diffusion Model.
Before the advent of DiT [77], the UNet architecture was predominantly used in diffusion models. We also conduct multi-task experiments based on a UNet pretrained model SDXL [78]. Specifically, we follow Marigold’s [47] approach by expanding the first convolution layer’s input channels from 4 to 8 to accommodate image inputs, and similarly use task prompts to guide the model in solving different tasks. However, as shown in Figure S16, we find that this approach is not sufficiently effective, even for a minimal multi-task scenario involving only depth and normal estimation. We attribute this limitation to the superior information aggregation capabilities of the Transformer architecture compared to UNet. While the UNet suffers from significant information loss during downsampling, the Transformer maintains more comprehensive representations, enabling it to perform better in multi-task scenarios.
5 Conclusion
We have introduced DiCeption, a multi-task visual generalist model based on the diffusion model. Our approach unifies different tasks in the RGB space, leveraging the prior knowledge of pre-trained image generation model to achieve results that are on par with or acceptably lower than the results of specialized foundation models. We achieve good performance without carefully cherry-picking extremely high-quality data or by using an exceptionally large amount of data.
Furthermore, for segmentation tasks, we demonstrate that the strategy of assigning random colors to different instances is highly effective in our framework, enabling high-quality entity segmentation and semantic segmentation. In few-shot fine-tuning, we are able to achieve high-quality results with minimal data and minimal trainable parameters. We believe that DiCeption sheds light on how to effectively use generative model priors to build a strong visual generalist, enabling more efficient solutions to perception tasks.
Appendix
Appendix A Dataset
Figure S1: Additional visualizations. Our one single model tackles multiple perception tasks.
We summarize the datasets used in our work in Table S1. The depth and normal data samples are obtained by randomly selecting 500K images from OpenImages [51] and labeling them using Depth Pro [7] and StableNormal [123], respectively. The 400K point segmentation data samples are obtained by randomly selecting images from the SA-1B dataset [49]. For the synthesis of point segmentation data, we extract the foreground from P3M-10K [55], AIM500 [57] and AM2K [56], randomly applying transformations such as rotation, resizing, and flipping. These transformed foregrounds are then pasted onto different background images, resulting in 200K synthetic images with fine-grained hair segmentation.
For the validation set, we evaluate depth using the same evaluation protocol as Genpercept [114], conducting tests on the NYUv2 [74], KITTI [32], ScanNet [23], DIODE [102], ETH3D [92]. Similarly, for normal estimation, we followed the evaluation protocol of StableNormal [123] and performed evaluations on the NYUv2 [74], ScanNet [23], DIODE [102]. For point segmentation, we conducted extensive comparisons across 23 datasets. The remaining tasks, including Entity Segmentation, Semantic Segmentation, and Human Keypoints, were evaluated on the MS COCO 2017 dataset [62]. We believe the comprehensive experiments provide solid evidence of the remarkable performance of our method.
Table S1: Dataset detail.
Training
Task Data Samples Dataset
Depth 500K OpenImages [51] + Depth Pro [7]
Normal 500K OpenImages [51] + StableNormal [123]
Point Segmentation 400K SA-1B [49]
Point Segmentation 200K P3M-10K [55], AIM500 [57] and AM2K [56]
Human Pose 42K MS COCO 2017 [62]
Semantic Segmentation 120K COCO-Rem [94]
Entity Segmentation 32K EntityV2 [81]
Validation
Task Dataset
Depth NYUv2 [74], KITTI [32], ScanNet [23], DIODE [102], ETH3D [92]
Normal NYUv2 [74], ScanNet [23], DIODE [102]
Point Segmentation PPDLS [71], DOORS [79], TimberSeg [30], NDD20 [101]
STREETS [95], iShape [120], ADE20K [134], OVIS [80]
Plittersdorf [37], EgoHOS [131], IBD [15], WoodScape [127]
TrashCan [41], GTEA [29, 60], NDISPark [20, 19], VISOR [24, 25]
LVIS [36], Hypersim [90], Cityscapes [22], DRAM [21]
BBBC038v1 [10], ZeroWaste [5], PIDRay [103]
Entity Segmentation MS COCO 2017 [62]
Semantic Segmentation MS COCO 2017 [62]
Human Keypoints MS COCO 2017 [62]
Appendix B Post-processing
Figure S2: Segmentation results on furry objects.
Algorithm 1 Keypoints Post-processing
0:  human pose RGB $\mathbf{x}$, GT keypoints $\mathbb{K}_{gt}$, RGB tolerance $\sigma$, distance threshold $\xi$
0:  extracted keypoints $\mathbb{K}_{pred}$
1:  $\mathbf{x}^{\prime}=\operatorname{ExtractRedRegions}(\mathbf{x},\,(255,0,0),\,\sigma)$
2:  $\mathbf{x}_{c}=\operatorname{GetConnectedComponents}(\mathbf{x}^{\prime})$
3:  $\mathbb{C}=\operatorname{GetCircular}(\mathbf{x}_{c})$
4:   $\mathbb{K}_{pred}=\varnothing$
5:  for $\mathbf{c}\in\mathbb{C}$ do
6:     $\mathbf{k}^{\prime}=\operatorname{ComputeCenterCoordinates}(\mathbf{c})$
7:     $d_{min}=\infty$
8:     for $\mathbf{k}\in\mathbb{K}_{gt}$ do
9:        $d=\operatorname{ComputeEuclideanDistance}(\mathbf{k}^{\prime},\mathbf{k})$
10:        if $d<d_{min}$  then
11:           $d_{min}=d$
12:           $t=\operatorname{GetKeypointType}(\mathbf{k})$
13:        end if
14:     end for
15:     if $d_{min}<\xi$ then
16:        continue
17:     end if
18:     $\mathbb{K}_{pred}=\mathbb{K}_{pred}\cup\{(\mathbf{k}^{\prime},t)\}$
19:  end for
20:  return  $\mathbb{K}_{pred}$
Algorithm 2 Segmentation Post-processing
0:  RGB segmentation mask $\mathbf{m}$, RGB tolerance $\sigma$, area threshold $\xi$, kernel size $k$, connected components number threshold $\eta$, duplicate mask threshold $\beta$
0:  extracted masks $\mathbb{M}_{pred}$
1:  Get the number of peaks $p$ of the histogram of $\mathbf{m}$
2:  Get the number of clusters $n=Mean(p)$
3:  Get the clustered colors by $\mathbb{C}=\operatorname{KMeans}(\mathbf{m},n)$
4:  $\mathbb{M}_{pred}=\varnothing$
5:  for $c\in\mathbb{C}$ do
6:     if $\operatorname{IsCloseToBlack(c,\sigma)}$  then
7:        continue
8:     end if
9:     $\mathbf{m}^{\prime}=\operatorname{GetMaskByRGB}(\mathbf{m},c,\sigma)$
10:     $\mathbf{m}^{\prime}=\operatorname{BinaryFillHoles}(\mathbf{m}^{\prime})$
11:     $\mathbf{m}^{\prime}=\operatorname{RefineWithMorphology}(\mathbf{m}^{\prime},k)$
12:     $a=\operatorname{GetArea}(\mathbf{m}^{\prime})$
13:     if $a<\xi$ then
14:        continue
15:     end if
16:     $y=\operatorname{GetConnectedComponentsNumber}(\mathbf{m}^{\prime})$
17:     if $y>\eta$ then
18:        continue
19:     end if
20:     $\mathbb{M}_{pred}=\mathbb{M}_{pred}\cup\{\mathbf{m}^{\prime}\}$
21:  end for
22:  $\mathbb{M}_{pred}=\operatorname{RemoveDuplicateMasks}(\mathbb{M}_{pred},\beta)$
23:  return  $\mathbb{M}_{pred}$
B.1 Post-processing for Keypoints
For keypoints, since all keypoints were labeled in red during training, our first step in post-processing is to extract all red regions from the RGB output. Next, we identify all connected components within the extracted red regions. For each connected component, we further extract sub-regions that approximate a circular shape. This step is crucial because, in some cases, multiple predicted keypoints may overlap, requiring us to separate them as much as possible. For example, when a person clasps his hands together, the keypoints for both hands may overlap.
Once the circular regions are identified, we compute their center points as the predicted keypoint coordinates. Since our model does not explicitly predict the type of each keypoint (e.g., hand, foot), we assign keypoint types by measuring the distance between the extracted keypoints and the ground truth (GT) keypoints. Each predicted keypoint is assigned the type of its nearest GT keypoint. To ensure robustness, we apply a distance threshold, considering only those predicted keypoints that are sufficiently close to a GT keypoint. Finally, all extracted keypoints that are successfully matched to a GT keypoint form our final predicted keypoint coordinates after post-processing. The algorithm is shown in Algorithm 1.
B.2 Post-processing for RGB Masks
For entity segmentation and semantic segmentation RGB masks, we employ clustering algorithms to extract the object masks. Specifically, we first compute the histogram peaks for each of the three RGB channels and estimate the number of clusters by averaging the peak counts across the three channels. We then use KMeans clustering to group the colors and identify the clustered regions in the RGB mask. For each identified cluster, we extract regions with RGB values close to the cluster’s centroid. This step is followed by morphological operations to refine the extracted masks, such as filling holes and removing small, fragmented regions. We further filter the masks by computing their area, excluding any regions that are too small to be meaningful.
Additionally, we also consider the number of connected components within the extracted masks, discarding overly fragmented results that have too many connected components. Finally, we refine the extracted masks by calculating the Intersection over Union (IoU) between them, removing any duplicate or overlapping masks. The algorithm is shown in Algorithm 2.
B.3 Performance Degradation of RGB Masks
Figure S3: When post-processing RGB masks, small regions and excessive numbers of objects lead to significant metric degradation.
Table S2: When post-processing RGB masks, small regions and excessive numbers of objects significantly lead to performance degradation.
Category AP $\uparrow$
Bear 76.3
Dog 68.9
Cat 71.7
Person 18.6
Bird 10.4
Book 10.8
We observe that while the quality of our semantic segmentation visualizations is high, the average precision (AP) for certain categories remains unsatisfactory. For example, for the Person category, we conducted exhaustive experiments and achieved good visualization results (highlighted by the green rectangle in Figure S3), but AP is low (as in Table S2). Although, clearly, there is room for us to improve the semantic segmentation results, we do not intend to fit the data bias of those existing datasets, as pointed out by other authors, e.g., [87].
We trace the root cause of this issue to degradation during post-processing, particularly due to small objects and an excessive number of objects. Specifically, during mask processing, we filter out small noise regions, but this also removes some positive samples, such as the crowd and the bird highlighted in red in rows 3 to 5 in Figure S3. However, if we do not filter these noise regions, they further degrade the results. In our setting, filtering noise regions results in better metrics compared to not filtering them. Additionally, when an image contains an excessive number of objects of the same category (as in row 6 of Figure S3), post-processing may erroneously group similarly colored but distinct objects into a single class, leading to lower metrics. Furthermore, as in Table S2, we examine categories with fewer small objects and instances of those categories, such as bear, dog, and cat, and observe higher AP scores. However, for categories with opposite characteristics, their AP scores tend to be lower.
Although we can optimize post-processing for individual datasets by adjusting hyperparameters for each image to achieve the best results, this approach becomes impractical for large-scale in-the-wild evaluation, as it requires significant manual effort. Consequently, the dependency on post-processing remains a limitation of our approach.
Figure S4: Additional few-shot fine-tuning results on image highlighting.
Appendix C Additional Results
C.1 Additional Visualizations
We present additional visualization results of our method across various tasks, as can be seen in Figures S1S2S4S6S7S8S9S10S11S12S13S14. For point-prompted segmentation, we further compared our approach with SAM. These results strongly demonstrate the potential of DiCeption. DiCeption is capable of achieving high-quality results, even in challenging scenarios. Furthermore, the few-shot fine-tuning of DiCeption, which requires minimal data and trainable parameters, strongly demonstrates the remarkable transferability of DiCeptionto tackle new tasks. Our DiCeption is capable of further refining the segmentation of fine details, such as intricate hair structures.
Figure S5: Our segmentation not only separates semantically identical objects but also distinguishes different instances of the same category, achieving higher segmentation quality. Moreover, One Diffusion tends to generate an image similar to the input when performing image understanding tasks, as red-highlighted in the figure.
Figure S6: Additional few-shot fine-tuning results on lung segmentation and tumor segmentation.
C.2 One-step Inference Does Not Work
Genpercept [114] demonstrates that one-step diffusion significantly enhances both the speed and accuracy of perceptual tasks. However, our experimental results reveal a notable increase in the proportion of failure cases when applying one-step diffusion in a multi-task setting, as illustrated in Figure S15. We believe that this is due to the potential overlap of denoising trajectories for different tasks. These overlapping trajectories can interfere with each other, resulting in failure cases. In contrast, multi-step denoising helps to mitigate such interactions. In a single-task setting, since the denoising trajectories pertain to a single task, this approach is more effective and stable.
Appendix D Limitations
Although our DiCeption achieves great results across multiple tasks, our model, as a diffusion model, leads to relatively longer inference times. On one H800, it takes an average of 0.8 seconds to process a single image. On one 4090-GPU card, inference for one image takes approximately 2 seconds. We believe that this issue can be addressed through few-step diffusion techniques. For tasks like semantic segmentation, the high demands of post-processing make the large-scale evaluation challenging. Despite the limitations, we believe that DiCeption is a valuable exploration for generalist visual foundation models.
Figure S7: Additional depth estimation visualizations.
Figure S8: Additional normal visualizations.
Figure S9: Additional entity segmentation visualizations.
Figure S10: Additional point-prompted segmentation visualizations.
Figure S11: Comparison of the segmentation results between DiCeption and SAM-vit-h with 1-point input.
Figure S12: Comparison of the segmentation results between DiCeption and SAM-vit-h with 5-point input.
Figure S13: Additional pose estimation visualizations.
Figure S14: Additional semantic segmentation visualizations.
Figure S15: The model tends to produce more failure cases in 1-step scenario.
Figure S16: A UNet-based model fails to perform multi-task.